two cards are drawn from a well shuffled ordinary deck of 52 cards. find the probability that they are both aces if the first card is (a) replaced (b) not replaced


1
Expert's answer
2021-06-14T15:08:32-0400

Replacement:

"P= \frac{4}{52} \times \frac{4}{52} = \frac{1}{169}"

Not replaced:

"P=\frac{4}{52} \times \frac{3}{51}= \frac{1}{221}"